OpenStudy (anonymous):

Note: Base 2= B2 logB2(y+2)-1=logB2(y-2)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

shay, I'll be right with you. Hope you've been able to solve that other problem. Have you?

OpenStudy (anonymous):

yes and thank you

OpenStudy (mathmale):

Shay, I'm back.

OpenStudy (mathmale):

logB2(y+2)-1=logB2(y-2) can be re-written as logB2(y+2)-logB2(y-2)=1. Please say whether you agree or disagree.

OpenStudy (anonymous):

yes, i agree

OpenStudy (anonymous):

don't you have to divide now?

OpenStudy (mathmale):

Use Rule #2 that I shared with you earliler: earlier, sorry Rule #2: log a-log b =log (a/b). can you now apply Rule #2 to combine the 2 terms on th left side of your equation?

OpenStudy (anonymous):

so it would be: logB2 y+2/y-2=1 ?

OpenStudy (mathmale):

That's pretty much correct, Shay, but it's important that you enclose y+2 and y-2 inside parentheses. Would you plese do that next time. But for now, write the equation 2 = 2

OpenStudy (mathmale):

and as the exponent of the first " 2 " please write logB2 [(y+2)/(y-2)]. (see how I used parentheses?

OpenStudy (mathmale):

and for the exponent of the 2nd " 2 " use 1. this new equation reduces to what? Note that we're using the inverse function property here, to simplify things.

OpenStudy (anonymous):

where did you get the 2 from? the one that's on the right side of the equal sign

OpenStudy (mathmale):

Yes, that's right. Perhaps you have a different way of doing this. if you find this confusing, I'll write everything out.

OpenStudy (anonymous):

so would it be 2^1= (y+2)(y-2)

OpenStudy (mathmale):

Here's what I'd expect to get as a result:\[2=\frac{ y+2 }{y-2 }\]

OpenStudy (anonymous):

would you multiply next or um, idk?

OpenStudy (mathmale):

the main difference is that y ou multiplied the ( )( ), whereas, I divided: ( ) / ( ) OK with this? then all you have to do is to solve the equation above for y. I'll be back in a few minutes in case you need to work on this further. so nice working with you!

OpenStudy (anonymous):

no i dont know what to do

OpenStudy (mathmale):

OK. If ... \[2 = \frac{ y+2 }{ y-2 }\] and we want to solve for y, one of the e3asier methods is to multiply both sides of the equation by the LCD, which is y-2. Then \[2(y-2) = y+2.\] Multiply out the left side. What do you get?

OpenStudy (mathmale):

I get 2y-4=y+2. How would you solve this equation for y?

OpenStudy (mathmale):

Hint #1: subtract y from both sides of the equation. Hint #2: add 4 to both sides of the equation. Solve for y. y = ?

OpenStudy (mathmale):

BRB

OpenStudy (mathmale):

Shay, If 2y-4=y+2, then subtracting y from both sides gives us y-4=2; adding 4 to both sides gives us y=2+4, or y=6. that's the solution. Please review this problem and determine whether or not you'd like to discuss any part of the solution in greater depth. BRB

OpenStudy (anonymous):

why did you out 2(y-2)=y+2 like how did you do that?

OpenStudy (mathmale):

Shay, I'm trying to solve the equation \[2=\frac{ y+2 }{ y-2 }\] for y. To do this, I multiply both sides of the equation by the denominator y-2, to eliminate the fraction: 2(y-2)=y+2. Are you OK with that?

OpenStudy (mathmale):

After multiplying out the first term, I get 2y-4=y+2 and then 2y-y-4=2 and then y - 4 = 2, and, finally, y=6.

OpenStudy (mathmale):

Shay, please note that if you substitute y=6 into \[2=\frac{ y+2 }{ y-2 }\]the resulting equation is true! This is the "solution" to the problem in logB2 that you presented.

OpenStudy (mathmale):

Hope you'll reply. If you do, I'll make sure we take this problem all the way to a satisfactory solution.